if f(x)=ax^2+2a and f(3)-f(1)=60, what is the value of a?
Answer by jim_thompson5910(35256)

f(x)=ax^2+2a

f(3)=a(3)^2+2a

f(3)=a*9+2a

f(3)=9a+2a

f(3)=11a

------------------------

f(x)=ax^2+2a

f(1)=a(1)^2+2a

f(1)=a*1+2a

f(1)=1a+2a

f(1)=3a

------------------------

f(3)-f(1)=60

11a - 3a = 60

8a = 60

a = 60/8

a = 15/2